OpenStudy (anonymous):

please help and i reward a medal For what value of x is sin x = cos 19°, where 0°< x < 90°? A.38° B.71° C.26° D.19°

OpenStudy (anonymous):

Take note of the following property: \[\sin \alpha=\cos 90-\alpha \] or the other way round. So\[x=90-19=71\] Answer, B.

OpenStudy (anonymous):

Like it @Hoslos, I'm stealing this solution!